OpenStudy (anonymous):

a rectangular fence is constructed that will enclose 100 ft sq of land. three sides will be built from fencing that costs 10 dollars per foot and the 4th side will be built from fencing that costs 23 dollars per foot. find the dimensions of the rectangular fence that minimizes the cost of materials. Please hurry!!!!!!!!!!!!!! I really need help fast!!

OpenStudy (anonymous):

use this as a guide A fence is to be built to enclose a rectangular area of 800 square feet. The fence along 3 sides is to be made of material $4 per foot. The material for the fourth side costs $12 per foot. Find the dimensions of the rectangle that will allow for the most economical fence to be built. if the expensive side is x and the other dimension is y, then the cost c is c = 4(x+2y) + 12x But, we know the area is xy=800, so y = 800/x and the cost is now c = 4(x+1600/x) + 12x minimum cost when dc/dx=0, so we need dc/dx = -16(400-x^2)/x^2 dc/dx=0 when x=20, so the fence is 20x40

OpenStudy (anonymous):

i will give u medal

OpenStudy (anonymous):

plzzz helpppp